OpenStudy (anonymous):

Assume that the area of the blue region is 100 units squared. Which equation represents this? A. x^2-15x+54=0 B. x^2-15x-46=0 C. x^2+15x+54=0 D. x^2+15x-46=0 I don't even know where to start. Can anyone explain this to me?

OpenStudy (anonymous):

Could we see the blue region first?

OpenStudy (anonymous):

that's the thing.... there is now blue region... there wasn't a picture

OpenStudy (anonymous):

I just took it as what the equation represents is 100 units squared, and im just really confused on that, how to find out which equation represents it I mean

OpenStudy (anonymous):

That's interesting. Which lesson is this problem for then? Using quadratic formula?

OpenStudy (anonymous):

No, I know how to use the quadratic formula, its and Apex highschool class and the lesson is on quadratic equation

OpenStudy (anonymous):

@agent0smith Is this even possible?

OpenStudy (anonymous):

@campbell_st

OpenStudy (agent0smith):

lol you can't answer this question without a pic. I'd guess it's either A or C, since both of those factor: A. x^2-15x+54= (x - 9)(x-6) C. x^2+15x+54= (x + 9)(x+6)

OpenStudy (anonymous):

apparently it was D, I got A or C too, but I was wrong I guess... It was Rigged, I swear

OpenStudy (agent0smith):

Well it's definitely rigged. You can't refer to a "blue region" and not have a diagram.

OpenStudy (anonymous):

That's What I thought!! I was like cool I can figure this out I think.... but then there wasn't a picture and I got totally messed up and confused, but whatever